Q2. Using the limiting law, Calculate the mean ionic activity coefficient at 25°C in water of 0.001 m of CaCl,.
Answer #1

Dear student, I have provided detailed solution for your question. Go through it. If you need further explanation feel free to comment.

Qa: Temperature 25°C = 298.15k limiting law: + lnf = -1.173/2, 2/JI = mean sonic actioľty coefficient I=908c strength chargeIn(14) = -1.173 ki ?/VE L-1.17 1 1732 72 m Q V.2 + V tt V z & 173 12 a Cu TV 2 0.00lm of cache dissociates as follows: cach (8 = expl-1.1731221 Q == 2 Cup tŲ z tt 0.001 expl-1.173|(+2)(+) (08]+[@@X-10] Q = exp op [1.23x2 pegel xo exp[- 2.346 * 0-4577